OpenStudy (anonymous):

Michelle draws a card from a standard deck of 52 cards. She replaces the card and draws a second card. What is the probability that she draws a red card and then a black card?

OpenStudy (anonymous):

First probability = 26 red cards/52 cards = 1/2 Second probability = 26 black cards/51 cards Total probability: 1/2*26/51 = 13/51

OpenStudy (anonymous):

that answer choice isn't on the test

OpenStudy (anonymous):

Opps, since she replaces the cards, it's 26 black cards/ 52 CARDS, which is 1/2*1/2 = 1/4 Sorry about that... too many questions

OpenStudy (anonymous):

thats okay .. and thanks alot !